Make-or-Buy, Traditional Analysis

Morrill Company produces two different types of gauges: a density gauge and a thickness gauge. The segmented income statement for a typical quarter follows.

Density
Gauge
Thickness
Gauge

Total
Sales $ 201,000 $ 107,200 $ 308,200
Less variable expenses 107,200 61,640 168,840
  Contribution margin $ 93,800 $ 45,560 $ 139,360
Less direct fixed expenses* 26,800 50,920 77,720
Segment margin $ 67,000 $ (5,360) $ 61,640
Less common fixed expenses 40,200
Operating income $ 21,440
* Includes depreciation.

The density gauge uses a subassembly that is purchased from an external supplier for $25 per unit. Each quarter, 2,680 subassemblies are purchased. All units produced are sold, and there are no ending inventories of subassemblies. Morrill is considering making the subassembly rather than buying it. Unit-level variable manufacturing costs are as follows:

Direct materials $2
Direct labor 3
Variable overhead 2

No significant non-unit-level costs are incurred.

Morrill is considering two alternatives to supply the productive capacity for the subassembly.

  1. Lease the needed space and equipment at a cost of $36,180 per quarter for the space and $13,400 per quarter for a supervisor. There are no other fixed expenses.
  2. Drop the thickness gauge. The equipment could be adapted with virtually no cost and the existing space utilized to produce the subassembly. The direct fixed expenses, including supervision, would be $50,920, $10,720 of which is depreciation on equipment. If the thickness gauge is dropped, sales of the density gauge will not be affected.

Required:

1. Should Morrill Company make or buy the subassembly?
Make the subassembly

If it makes the subassembly, which alternative should be chosen?
Drop the thickness gauge

Enter the relevant costs of each alternative.

Lease and Make Buy Drop Thickness Gauge and Make
Total relevant costs $ $ $

2. Suppose that dropping the thickness gauge will decrease sales of the density gauge by 10 percent. What decision should now be made?

3. Assume that dropping the thickness gauge decreases sales of the density gauge by 10 percent and that 3,752 subassemblies are required per quarter. As before, assume that there are no ending inventories of subassemblies and that all units produced are sold. Assume also that the per-unit sales price and variable costs are the same as in Requirement 1. Include the leasing alternative in your consideration. Now, what is the correct decision?
Answer #1

1.a. Let's consider first option ie. Buy the subassembly.

Total Cost = 25*2680 = $ 67,000

b. Second option: Lease the place and make subassembly:
Total Cost = Fixed expenses + Variable expenses = (Rent+Supervisor salary) + Variable cost

= 36180 + 13400 + (2+3+2)*2680 = 49580 + 7*2680 =$ 68,340

c. third option: drop the thickness gauge and make subassembly

Total cost = Fixed expenses + Variable expenses = 50920 + 7*2680 = $ 69,680

2. if density gauge sales are dropped by 10% with no anticipation, then total cost should factor in the opportunity cost of this decision.

Total cost = 58960 + (10% of 201000) = 69680+20100 = $ 89780

3.Now further, 3752 subassemblies are required and 10% sales are dropped.

Total cost = Fixed expenses + Variable expenses + (10% decrement in sales) = 50920-10720 + 7*3752 + 10%*201000 = $86564